# Possible Causes of False General Relativity Violations in Gravitational Wave Observations

### v1

@@ 3.1.3 Kick-induced Effects
The anisotropic emission of GWs during a CBC carries away linear momentum and results in a recoil or kick of the merger remnant (Fitchett, 1983; Favata et al., 2004). The kick leaves the following imprints in the GW signal: the Doppler effect (Favata, 2009a) and the aberration effect (Torres-Orjuela et al., 2020) on the post-merger signal along with an additional contribution of a (linear) memory effect (Favata, 2009a) to the whole GW signal (Mahapatra et al., 2023). Since the black hole kicks are non-relativistic, the kick-induced effects are small and might not be important for current GW detectors but could be crucial for XG detectors (Gerosa & Moore, 2016; Mahapatra et al., 2023). For loud ringdown signals (SNR100greater-than-or-equivalent-toabsent100$\gtrsim$ 100≳ 100,  Gerosa & Moore (2016)) in the XG era, these kick-induced effects, if not accounted appropriately in the waveform model (Boyle, 2016; Varma et al., 2019), might contaminate those tests of GR that depend on the post-merger signal and kick (Varma et al., 2020) of the remnant (see, e.g., Hughes & Menou (2005); Ghosh et al. (2016); Mahapatra et al. (2023); Isi et al. (2019b); Carullo et al. (2019); Maggio et al. (2023)).
@@ 3.1.4 Beyond Fundamental Modes in Ringdown Signal
The gravitational radiation from a perturbed black hole is in the form of quasi-normal modes (Vishveshwara, 1970; Press, 1971). At sufficiently late times following a binary black hole merger, it is expected that the remnant can be very well approximated by a perturbed Kerr black hole. Moreover, it is well known that the radiation at this stage is dominated by just the fundamental quasi-normal mode, since it is the slowest damped quasi-normal-mode (QNM) (Detweiler, 1980; Dreyer et al., 2004; Berti et al., 2006). The frequency and damping time of a mode are in one-to-one correspondence with the remnant mass and spin. In principle, assuming GR and using NR simulations, the latter quantities could be predicted from the properties of the progenitor binary, which can be extracted from the premerger signal. In practice, waveform systematics in the premerger phase could jeopardize this ringdown consistency test (Dhani et al., 2024). For example, large unmodelled eccentricity could lead to an inconsistency in the final mass and spin, and hence to a false GR deviation (Narayan et al., 2023). In the spirit of the original black-hole spectroscopy program (Detweiler, 1980; Dreyer et al., 2004; Berti et al., 2006; Gossan et al., 2012), it is therefore better to test GR using ringdown signals only, and an “agnostic” selection of multiple modes to model the ringdown (Baibhav et al., 2023).
@@
Recently, there have been efforts to increase the range of validity of linear perturbation theory by modeling the early postmerger signal using overtones and mirror modes (Baibhav et al., 2018; Baibhav & Berti, 2019; Giesler et al., 2019; Dhani, 2021; Dhani & Sathyaprakash, 2021; Magaña Zertuche et al., 2022; Ma et al., 2022, 2023a; Baibhav et al., 2023; Cheung et al., 2024; Takahashi & Motohashi, 2023; Qiu et al., 2024; Clarke et al., 2024). These studies show that the inclusion of these additional QNMs improve the remnant mass and spin estimates using a ringdown model. They also show that there will be biases in the remnant parameters if a ringdown model is used to describe early postmerger without the inclusion of such QNMs. Such biases in parameter estimation can show a deviation from the predictions of GR. Isi & Farr (2021) investigated the impact of an incomplete ringdown model on parameter recovery by analyzing a synthetic signal mimicking a binary black hole ringdown (see also Baibhav et al. (2023) for a discussion). Their findings reveal biased parameter measurements in instances of very high ringdown SNR. Dhani & Sathyaprakash (2021) displayed the modulations in the odd-$m$ modes in the waveform and how the inclusion of mirror modes in the ringdown waveform model can explain these modulations.
@@
There are claims in the literature that overtones have been detected (Isi & Farr, 2022; Finch & Moore, 2022; Ma et al., 2023b) and used to test the “no-hair” theorem with GW150914 (Isi et al., 2019b). However, there is a disagreement in the literature regarding the significance of the measurement of the first overtone in GW150914 (Finch & Moore, 2022; Cotesta et al., 2022; Crisostomi et al., 2023; Gennari et al., 2024; Correia et al., 2023). There are also theoretical arguments suggesting caution in the use of overtones for no-hair theorem tests (Sberna et al., 2022a; Lagos & Hui, 2023; Baibhav et al., 2023; Nee et al., 2023; Zhu et al., 2024a; Cheung et al., 2024). The above authors show, using toy models, black hole perturbation theory and NR simulations, that even though the estimates of the final mass and spin of the black hole can be improved starting the ringdown analysis at earlier times by the addition of overtones, a linear model including only overtones is not appropriate at early times (see also Bhagwat et al. (2020)). Therefore, they contend that overtones are unphysical and that their role in a waveform model is to “fit away” other features in the signal, namely, transients related to the initial data, power-law tails at late times, and nonlinearities.
@@
However, for less symmetric binaries than GW150914 (as commonly expected among current and future catalogs) the original black-hole spectroscopy program can be realized using higher-order modes in addition to the least damped QNM, i.e., $(l,|m|)=(3,3),(2,1),(4,4)$, can be used to perform independent tests of the no-hair theorem (Berti et al., 2007a, 2016; Brito et al., 2018; Carullo et al., 2019; Cabero et al., 2020; Jiménez Forteza et al., 2020; Ghosh et al., 2021; Gennari et al., 2024; Bhagwat et al., 2023). Given current estimates of the merger rates, XG detectors are predicted to perform percent-accuracy tests for a few events per year (Berti et al., 2016; Ota & Chirenti, 2020, 2022; Bhagwat et al., 2023).
@@
To conduct any of the above tests of GR using the perturbative ringdown model, one must make a choice on the start time of the ringdown to begin fitting exponentially damped sinusoids. The analysis should begin as soon as the perturbative prescription is relevant. On one hand, waiting too long to begin the analysis will make testing GR impossible because the strain amplitude has decayed exponentially (e.g., Thrane et al., 2017; Bustillo et al., 2021). However, beginning the analysis too early could result in overfitting to non-linear features in the signal (e.g., Bhagwat et al., 2020; Baibhav et al., 2023). To undertake robust tests of GR, some criterion for the analysis start time should be established through, e.g., searching for the earliest time at which one can measure self-consistent QNM parameters with time (Cheung et al., 2024; Takahashi & Motohashi, 2023; Clarke et al., 2024). A further source of systematics is the decomposition of QNMs in spherical rather than spheroidal harmonics; if unmodelled, the spherical-spheroidal mode mixing introduces biases for highly spinning remnants (Baibhav et al., 2023).
@@
Another important effect of the nonlinearity in the ringdown stage is the presence of second-order QNMs (Mitman et al., 2023; Cheung et al., 2023; Khera et al., 2023), which are generated through mode-mode couplings. The frequency of a second-order QNM is twice as the associated “parent” linear QNM. Its amplitude and phase are also uniquely determined by the linear mode (Zhu et al., 2024b; Redondo-Yuste et al., 2023; Ma & Yang, 2024), as a nontrivial prediction of GR at the nonlinear level. The dominant nonlinear modes may be observable with XG detectors, although event rates are uncertain (Yi et al., 2024).
@@
An approach complementary to null tests using QNM frequencies and damping times is to test QNM amplitude-phase relations predicted by NR simulations within GR. This test was successfully applied to GW190521 in Forteza et al. (2023), finding that measurement errors for this event are still large, but would strongly improve for the louder detections routinely expected for XG detectors.
@@
Finally, because of its short duration, one should be careful with the statistical methods and their underlying assumptions while analyzing the ringdown signal. Seemingly innocuous data processing choices such as the uncertain starting time, duration of the signal, and noise estimation techniques can lead to materially different inferences (Isi et al., 2019b; Cotesta et al., 2022; Isi & Farr, 2023; Carullo et al., 2023; Crisostomi et al., 2023; Wang et al., 2023). While the ringdown signal is typically analyzed in the time domain, frequency domain methods have also been proposed (Capano et al., 2023; Finch & Moore, 2021; Ma et al., 2022; Crisostomi et al., 2023) with the approach of Capano et al. (2023) shown to be formally equivalent to the time-domain approach (Isi & Farr, 2021). Even then, Capano et al. (2023) comes to a different conclusion regarding the ringdown of GW190521 compared to Abbott et al. (2021a) or Siegel et al. (2023). This highlights the need to better understand systematics and data analysis techniques in the analysis of ringdown signals.
@@ 3.2.1 Higher-order Modes, Precession, and Memory
Gravitational waveforms can be decomposed in the basis of spin weighted spherical harmonics with spin weight $s=-2$, $Y^{lm}_{-2}(\iota)$, where $\iota$ is the inclination angle. In this basis, for nonprecessing systems, the dominant contribution to the GW amplitude comes from the $(l,|m|)=(2,2)$ harmonics. The $(2,1)$ and $(3,3)$ harmonics are subdominant and suppressed by a prefactor that goes to 00 for symmetric (equal mass) binaries (Berti et al., 2007b; Kidder, 2008; Arun et al., 2009; Pan et al., 2011; Blanchet, 2014). These modes only contribute for systems that are not face-on/off ($\iota\neq 0,2\pi$), and become particularly important for unequal mass binaries. The presence of these higher-order modes causes characteristic modulations in the amplitude and phase of the waveform.
@@
The effect of higher-order modes becomes even more important in the presence of spin-induced precession. Spin-induced precession occurs when the spin angular momentum vectors of the binary components are not aligned with the orbital angular momentum vector, leading to the precession of the orbital angular momentum (or, equivalently, the orbital plane of the binary) as well as the spin vectors about the total angular momentum of the binary. The effect of precession is best understood by considering two frames of reference (Buonanno et al., 2003; Schmidt et al., 2012; O’Shaughnessy et al., 2013)—the inertial frame in which the binary appears to be precessing, and the co-precessing frame that follows the instantaneous motion of the orbital plane where the effects of precession disappear. The inertial modes can then be approximately described as the sum of nonprecessing modes with the same $l$ value and all possible $m$ values, each rotated using Wigner D-matrices which depend on the instantaneous position of the orbital plane (Hannam et al., 2014). Thus, due to spin-induced precession, subdominant precessing modes will have contributions from both dominant and subdominant nonprecessing modes, increasing the precession effect due to the presence of higher-order modes in the waveform (Islam et al., 2021).
@@
A consequence of using nonprecessing modes to approximate the co-precessing-frame signal is that these obey the reflection symmetry $h_{\ell m}=(-1)^{\ell}h_{\ell-m}^{*}$, which no longer holds for precessing binaries  (Boyle et al., 2014; Ramos-Buades et al., 2020). Most state-of-the-art waveform models, with the exception of NRSur7dq4 (Varma et al., 2019) and IMRPhenomXO4a (Thompson et al., 2023; Ghosh et al., 2024), currently rely on this approximation. While the impact of anti-symmetric contributions to the waveform modes is typically small, neglecting these effects could result in biased measurements of the spin magnitude and orientation at high SNR (Kalaghatgi & Hannam, 2021; Kolitsidou et al., 2024).
@@
Currently, state-of-the-art nonprecessing waveforms like IMRPhenomXHM (García-Quirós et al., 2020) include the harmonics $(l,|m|)=(2,1),(3,3),(3,2),(4,4)$, and SEOBNRv5HM (Pompili et al., 2023), in addition to these, also includes $(l,|m|)=(4,3)$ and $(5,5)$. Their precessing counterparts are IMRPhenomXPHM (Pratten et al., 2021) and SEOBNRv5PHM (Ramos-Buades et al., 2023a), respectively. The widely used NR surrogate waveform model, NRSur7dq4, has been trained with simulations with mass ratio less than 4, and contains all spherical-harmonic modes with $l\leq 4$.
@@
Many studies have explored the improvement in the inference of source parameters due to the inclusion of spin-induced orbital precession and higher-order modes (Klein et al., 2009; Shaik et al., 2020; Krishnendu & Ohme, 2022; Loutrel et al., 2023). Particularly, for edge-on systems, including higher-order modes improves parameter estimation by breaking the luminosity distance-inclination angle degeneracy, whereas modulations due to spin-induced precession break the degeneracy between the spin and mass parameters. Additionally, the amplitude of the higher-order modes also brings information about the mass ratio of the source.
@@
We should note that none of these models discussed above contain the memory modes that depend on the binary’s past history. The most well-known of these is the displacement memory effect which is dominant in the $l=2,m=0$ mode, and the next leading memory effect, known as the spin memory, is dominant in $l=3,m=0$ mode for the non-precessing binaries (see e.g., Favata (2009b), and Nichols (2017)). There are other higher-order memory effects, but these can be extremely sub-dominant. Most of these are discussed in Mitman et al. (2020) and references therein. While these are small effects, they will need to be included to prevent biases, and have now been included in a surrogate model for nonprecessing (quasicircular) binary black holes constructed using the waveforms obtained from Cauchy-characteristic evolution  (Yoo et al., 2023). The effect of non-linear memory on the binary black hole parameter estimation is studied in Xu et al. (2024) where the dominant displacement memory in the $l=2,m=0$ mode starts to affect the parameter inference at ${\rm SNR}>60$ for the current generation ground-based detectors (such as LIGO A#). Moreover, the effect of memory has been studied in the case of neutron star-black hole and binary neutron star mergers (Tiwari et al., 2021; Lopez et al., 2024), where it is argued that the memory can affect parameter estimation for the XG detectors.
@@
Therefore, analyzing a GW signal that has a significant magnitude of spin-induced precession, higher order mode content, and memory effect with an inaccurate or incomplete waveform model may not only deteriorate parameter estimation, but also show biases in the inference of other source parameters (see, e.g., Islam et al. (2021)). A recent study has investigated systematics due to waveform mismodeling by comparing SEOBNRv5PHM and IMRPhenomXPHM. It was found that systematic biases can impact the current and future GW-detector networks, affecting the inference of realistic binary black hole population properties, as well as, the science cases of individual loud signals (Dhani et al., 2024), and more in general binaries with large mass ratios and high precession. Such systematic biases may eventually find their way into the measurement of a beyond-GR parameter depending on the nature of its correlation with the other source parameters, inducing a false violation of GR. Hence, it is essential to use accurate waveform models with spin-precession effects, sufficient number of higher-order modes, and memory effects while testing a GW signal for a violation of GR.

### v2

@@ 3.1.3 Kick-induced Effects
The anisotropic emission of GWs during a CBC carries away linear momentum and results in a recoil or kick of the merger remnant [230, 231]. The kick leaves the following imprints in the GW signal: the Doppler effect [232] and the aberration effect [233] on the post-merger signal along with an additional contribution of a (linear) memory effect [232] to the whole GW signal [234]. Since the black hole kicks are non-relativistic, the kick-induced effects are small and might not be important for current GW detectors but could be crucial for XG detectors [235, 234]. For loud ringdown signals (SNR100greater-than-or-equivalent-toabsent100$\gtrsim$ 100≳ 100,  [235]) in the XG era, these kick-induced effects, if not accounted appropriately in the waveform model [236, 237], might contaminate those tests of GR that depend on the post-merger signal and kick [238] of the remnant (see, e.g., [239, 240, 234, 241, 242, 39]).
@@ 3.1.4 Beyond Fundamental Modes in Ringdown Signal
The gravitational radiation from a perturbed black hole is in the form of quasi-normal modes [243, 244]. At sufficiently late times following a binary black hole merger, it is expected that the remnant can be very well approximated by a perturbed Kerr black hole. Moreover, it is well known that the radiation at this stage is dominated by just the fundamental quasi-normal mode, since it is the slowest damped quasi-normal-mode (QNM) [245, 246, 247]. The frequency and damping time of a mode are in one-to-one correspondence with the remnant mass and spin. In principle, assuming GR and using NR simulations, the latter quantities could be predicted from the properties of the progenitor binary, which can be extracted from the premerger signal. In practice, waveform systematics in the premerger phase could jeopardize this ringdown consistency test [248]. For example, large unmodelled eccentricity could lead to an inconsistency in the final mass and spin, and hence to a false GR deviation [135]. In the spirit of the original black-hole spectroscopy program [245, 246, 247, 249], it is therefore better to test GR using ringdown signals only, and an “agnostic” selection of multiple modes to model the ringdown [250].
@@
Recently, there have been efforts to increase the range of validity of linear perturbation theory by modeling the early postmerger signal using overtones and mirror modes [251, 252, 253, 254, 255, 256, 257, 258, 250, 259, 260, 261, 262]. These studies show that the inclusion of these additional QNMs improve the remnant mass and spin estimates using a ringdown model. They also show that there will be biases in the remnant parameters if a ringdown model is used to describe early postmerger without the inclusion of such QNMs. Such biases in parameter estimation can show a deviation from the predictions of GR. Isi & Farr [263] investigated the impact of an incomplete ringdown model on parameter recovery by analyzing a synthetic signal mimicking a binary black hole ringdown (see also [250] for a discussion). Their findings reveal biased parameter measurements in instances of very high ringdown SNR. Dhani & Sathyaprakash [255] displayed the modulations in the odd-$m$ modes in the waveform and how the inclusion of mirror modes in the ringdown waveform model can explain these modulations.
@@
The black hole spectroscopy program in GW literature aims to test the Kerr nature from the observed QNM spectrum. These tests are typically referred to as “no-hair” theorem tests too. However, since the tests are based on QNMs as the only observables, they are not sensitive to the type of black hole hair — namely, primary hair [^1] In this context, primary hair refers to extra charges that are independent of the black hole mass and spin (e.g., the electric charge in the Kerr-Newman solution), whereas secondary hair refers to extra charges that are fixed in terms of the mass and spin. or secondary hair [264]. Therefore, any modification to the Kerr QNM spectra would fall under these tests. There are claims in the literature that overtones have been detected [265, 266, 267] and used to test the no-hair theorem with GW150914 [241]. However, there is a disagreement in the literature regarding the significance of the measurement of the first overtone in GW150914 [266, 268, 269, 270, 271]. There are also theoretical arguments suggesting caution in the use of overtones for no-hair theorem tests [272, 273, 250, 274, 275, 259]. The above authors show, using toy models, black hole perturbation theory and NR simulations, that even though the estimates of the final mass and spin of the black hole can be improved starting the ringdown analysis at earlier times by the addition of overtones, a linear model including only overtones is not appropriate at early times (see also [276]). Therefore, they contend that overtones are unphysical and that their role in a waveform model is to “fit away” other features in the signal, namely, transients related to the initial data, power-law tails at late times, and nonlinearities.
@@
However, for less symmetric binaries than GW150914 (as commonly expected among current and future catalogs) the original black-hole spectroscopy program can be realized using higher-order modes in addition to the least damped QNM, i.e., $(l,|m|)=(3,3),(2,1),(4,4)$, can be used to perform independent tests of the no-hair theorem [277, 278, 279, 242, 280, 281, 282, 270, 283]. Given current estimates of the merger rates, XG detectors are predicted to perform percent-accuracy tests for a few events per year [278, 284, 285, 283].
@@
To conduct any of the above tests of GR using the perturbative ringdown model, one must make a choice on the start time of the ringdown to begin fitting exponentially damped sinusoids. The analysis should begin as soon as the perturbative prescription is relevant. On one hand, waiting too long to begin the analysis will make testing GR impossible because the strain amplitude has decayed exponentially (e.g., [286, 287]). However, beginning the analysis too early could result in overfitting to non-linear features in the signal (e.g., [288, 250]). To undertake robust tests of GR, some criterion for the analysis start time should be established through, e.g., searching for the earliest time at which one can measure self-consistent QNM parameters with time [259, 260, 262]. A further source of systematics is the decomposition of QNMs in spherical rather than spheroidal harmonics; if unmodelled, the spherical-spheroidal mode mixing introduces biases for highly spinning remnants [250].
@@
Another important effect of the nonlinearity in the ringdown stage is the presence of second-order QNMs [289, 290, 291], which are generated through mode-mode couplings. The frequency of a second-order QNM is twice as the associated “parent” linear QNM. Its amplitude and phase are also uniquely determined by the linear mode [292, 293, 294], as a nontrivial prediction of GR at the nonlinear level. The dominant nonlinear modes may be observable with XG detectors, although event rates are uncertain [295].
@@
An approach complementary to null tests using QNM frequencies and damping times is to test QNM amplitude-phase relations predicted by NR simulations within GR. This test was successfully applied to GW190521 in [296], finding that measurement errors for this event are still large, but would strongly improve for the louder detections routinely expected for XG detectors.
@@
Finally, because of its short duration, one should be careful with the statistical methods and their underlying assumptions while analyzing the ringdown signal. Seemingly innocuous data processing choices such as the uncertain starting time, duration of the signal, and noise estimation techniques can lead to materially different inferences [241, 268, 297, 298, 269, 299]. While the ringdown signal is typically analyzed in the time domain, frequency domain methods have also been proposed [300, 301, 257, 269] with the approach of [300] shown to be formally equivalent to the time-domain approach [263]. Even then, [300] comes to a different conclusion regarding the ringdown of GW190521 compared to [4] or [302]. This highlights the need to better understand systematics and data analysis techniques in the analysis of ringdown signals.
@@ 3.2.1 Higher-order Modes, Precession, and Memory
Gravitational waveforms can be decomposed in the basis of spin weighted spherical harmonics with spin weight $s=-2$, $Y^{lm}_{-2}(\iota)$, where $\iota$ is the inclination angle. In this basis, for nonprecessing systems, the dominant contribution to the GW amplitude comes from the $(l,|m|)=(2,2)$ harmonics. The $(2,1)$ and $(3,3)$ harmonics are subdominant and suppressed by a prefactor that goes to 00 for symmetric (equal mass) binaries [303, 304, 305, 306, 307]. These modes only contribute for systems that are not face-on/off ($\iota\neq 0,2\pi$), and become particularly important for unequal mass binaries. The presence of these higher-order modes causes characteristic modulations in the amplitude and phase of the waveform.
@@
The effect of higher-order modes becomes even more important in the presence of spin-induced precession. Spin-induced precession occurs when the spin angular momentum vectors of the binary components are not aligned with the orbital angular momentum vector, leading to the precession of the orbital angular momentum (or, equivalently, the orbital plane of the binary) as well as the spin vectors about the total angular momentum of the binary. The effect of precession is best understood by considering two frames of reference [308, 309, 310]—the inertial frame in which the binary appears to be precessing, and the co-precessing frame that follows the instantaneous motion of the orbital plane where the effects of precession disappear. The inertial modes can then be approximately described as the sum of nonprecessing modes with the same $l$ value and all possible $m$ values, each rotated using Wigner D-matrices which depend on the instantaneous position of the orbital plane [311]. Thus, due to spin-induced precession, subdominant precessing modes will have contributions from both dominant and subdominant nonprecessing modes, increasing the precession effect due to the presence of higher-order modes in the waveform [312].
@@
A consequence of using nonprecessing modes to approximate the co-precessing-frame signal is that these obey the reflection symmetry $h_{\ell m}=(-1)^{\ell}h_{\ell-m}^{*}$, which no longer holds for precessing binaries  [313, 314]. Most state-of-the-art waveform models, with the exception of NRSur7dq4 [237] and IMRPhenomXO4a [315, 316], currently rely on this approximation. While the impact of anti-symmetric contributions to the waveform modes is typically small, neglecting these effects could result in biased measurements of the spin magnitude and orientation at high SNR [317, 318].
@@
Currently, state-of-the-art nonprecessing waveforms like IMRPhenomXHM [319] include the harmonics $(l,|m|)=(2,1),(3,3),(3,2),(4,4)$, and SEOBNRv5HM [320], in addition to these, also includes $(l,|m|)=(4,3)$ and $(5,5)$. Their precessing counterparts are IMRPhenomXPHM [321] and SEOBNRv5PHM [322], respectively. The widely used NR surrogate waveform model, NRSur7dq4, has been trained with simulations with mass ratio less than 4, and contains all spherical-harmonic modes with $l\leq 4$.
@@
Many studies have explored the improvement in the inference of source parameters due to the inclusion of spin-induced orbital precession and higher-order modes [323, 324, 325, 326]. Particularly, for edge-on systems, including higher-order modes improves parameter estimation by breaking the luminosity distance-inclination angle degeneracy, whereas modulations due to spin-induced precession break the degeneracy between the spin and mass parameters. Additionally, the amplitude of the higher-order modes also brings information about the mass ratio of the source.
@@
We should note that none of these models discussed above contain the memory modes that depend on the binary’s past history. The most well-known of these is the displacement memory effect which is dominant in the $l=2,m=0$ mode, and the next leading memory effect, known as the spin memory, is dominant in $l=3,m=0$ mode for the non-precessing binaries (see e.g., [327] and [328]). There are other higher-order memory effects, but these can be extremely sub-dominant. Most of these are discussed in [329] and references therein. While these are small effects, they will need to be included to prevent biases, and have now been included in a surrogate model for nonprecessing (quasicircular) binary black holes constructed using the waveforms obtained from Cauchy-characteristic evolution  [330]. The effect of non-linear memory on the binary black hole parameter estimation is studied in [331] where the dominant displacement memory in the $l=2,m=0$ mode starts to affect the parameter inference at ${\rm SNR}>90$ for the current generation ground-based detectors (such as LIGO A#). Moreover, the effect of memory has been studied in the case of neutron star-black hole and binary neutron star mergers [332, 333], where it is argued that the memory can affect parameter estimation for the XG detectors. Studies show that it will be difficult to detect the presence of memory in individual sources with the current LIGO, Virgo, and KAGRA detectors at O4 sensitivity or even O5 sensitivity, but it could be detected in a population using the stacking procedure (e.g., [334]). Thus, it is necessary to understand the effect of memory on parameter estimation and tests of GR at the population-level.
@@
Therefore, analyzing a GW signal that has a significant magnitude of spin-induced precession, higher order mode content, and memory effect with an inaccurate or incomplete waveform model may not only deteriorate parameter estimation, but also show biases in the inference of other source parameters (see, e.g., [312]). A recent study has investigated systematics due to waveform mismodeling by comparing SEOBNRv5PHM and IMRPhenomXPHM. It was found that systematic biases can impact the current and future GW-detector networks, affecting the inference of realistic binary black hole population properties, as well as, the science cases of individual loud signals [248], and more in general binaries with large mass ratios and high precession. Such systematic biases may eventually find their way into the measurement of a beyond-GR parameter depending on the nature of its correlation with the other source parameters, inducing a false violation of GR. Hence, it is essential to use accurate waveform models with spin-precession effects, sufficient number of higher-order modes, and memory effects while testing a GW signal for a violation of GR.